Laser Waves with wavelength 633 nm have 3 Mw power. What will be the pressure applied by these waves at 3 mm^2 Cross section area.

( Given that Surface is 100% reflecting ( R =1) and Ray's fall perpendicularly to the surface)
Answer #1

Intensity, I= 2 = 3 x 103 W = 1000 W/m² 3x10-6 m² ... Radiation Pressure : -6 N/m² = 6-67 x 100 Awower =